# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ IV-C Learning-based Contact Planning
Conclusion: While significant progress has been made in humanoid multi-contact planning, future work should focus on developing more integrated approaches that combine the strengths of search-based, optimization-based, and learning-based methods. Specifically, addressing the computational complexity of CIP and improving real-time performance will be key. Future directions could explore hybrid approaches that incorporate efficient contact sequence generation/contact dynamics, apply contact-implicit constraints in real-time, and achieve learning-based contact prediction to enhance robustness and adaptability in complex loco-manipulation tasks. The readers are recommended to further read the survey on humanoid multi-contact planning [141].
@@ V-A Simplified Models
TABLE II: Recent MPC Approaches on Humanoid Loco-manipulation
Paper Robot Model Interaction Modeling Method Locomotion (L) and Manipulation (M) Planning MPC Frequency Solving Method
[185] SRBM Optimizing external wrench(es) at contact(s) Unified $20$ Hz QP
[2] SRBM Predefined external force Unified $300$ Hz QP
[186] SRBM Negligible object dynamics Separated: L: MPC; M: Keyframe interpolation -- QP
[187] LIPM Negligible object dynamics Separated: L: MPC; M: Retargeting through teleoperation -- QP
[117] CD Optimizing external wrench(es) at contact(s) Unified offline SQP
[28] CD Predefined external wrench Unified -- QP
[188] CD Optimizing external wrench(es) at contact(s) Unified $5$ Hz/offline SQP
[189] CD Optimizing external wrench(es) at contact(s) Unified $10$ Hz Interior-point
[183] CD Estimated as external wrench through sensors Unified $5$ Hz Interior-point
[190] WBD Optimizing external wrench(es) at contact(s) Unified $100$ Hz DDP
@@
On the other hand, the Linear Inverted Pendulum Model (LIPM) has been a popular choice as a linearized dynamics model for humanoid locomotion [23] and in multi-contact [191]. To further extend the capability of rough terrain walking, the Prismatic Inverted Pendulum Model (PIPM) is used to allow CoM movements on a non-flat surface [192, 193]. Extending LIPM to loco-manipulation tasks is achieved through teleoperation [187]. However, such model inherently lacks the ability to address contact interactions and loco-manipulation dynamics, necessitating a lower-level whole-body control for balancing and manipulation tasks.
@@ V-B Whole-body Models
While simplified dynamics models offer computation efficiency, they often lack model accuracy and have limited capability of whole-body motion planning due to high simplification assumptions. Conversely, whole-body models are more accurate representations of the robot dynamics and are better suited for planning versatile motions and interactions with objects and the environment. NMPC comes into play when constraints or cost functions become nonlinear, for example, dynamics constraints formed by kinodynamics and Whole-Body Dynamics (WBD).
@@
On the other hand, leveraging the WBD in MPC has become a popular approach recently. Joint-space WBD, described in (5), accurately models a free-floating articulated robot such as a humanoid robot. WBD provides flexibility in defining arbitrary and allowable contact in dynamics modeling, including interaction forces through manipulation. However, the inherent high nonlinearity and nonconvexity impose significant computational burdens on WBD-based Nonlinear Programs (NLP), making them challenging for real-time planning. The numerical accuracy of solutions is often compromised to exchange for real-time application (e.g.[6]), even with accelerated solving methods for NLPs (e.g., Sequential Quadratic Programming (SQP), Differential Dynamic Programming (DDP)). This computation issue is pronounced in high-degree-of-freedom humanoids performing loco-manipulation tasks, such as human-robot payload transportation, that demand additional modeling of object dynamics, safety-critical control, and robust trajectory solutions. Therefore, a major focus of this section is the discussion of different approaches to speed up NMPC while maintaining its solution accuracy.
@@ V-C Mixed-fidelity Models
As one way to mix models, cascaded-fidelity models (a.k.a. hierarchical dynamics) use different models to govern different segments of the horizon [195, 196, 197]. These methods typically employ high-fidelity (e.g., full-order) models in near horizons and low-fidelity (e.g., simple) models for distant horizons, thus maintaining the solution accuracy in near horizons while solving the myopic issue by allowing a longer horizon using simple models. This approach could be suitable in loco-manipulation tasks as it would either simplify interaction dynamics as simple external forces or impose the object dynamics as part of CD in the far horizons to allow improved real-time computation compared to full dynamics models.
Another way to mix models is to have different dynamics models overlap between their horizons. In such cases, achieving a consensus between these overlapped models is necessary. To solve problems with such mixed-fidelity models, [198] decomposes a single TO that incorporates both dynamics into two subproblems and then alternates between the two to achieve consensus. In a similar vein, [199] alternate between the CD and WBK subproblems. Overall, model simplification over MPC horizons will remain an effective approach [7]. On the other hand, mixed-fidelity models demonstrate superior capability but require careful consideration of combined models.
@@ V-D NMPC Speed-up
NMPC Speed-up via Structure Exploitation: NMPC problems often involve complex dynamics and constraints that can be computationally intensive to solve. Exploiting the structure within these problems can significantly enhance their solvability and efficiency, such as extracting variables that directly interact with each other, identifying repetitive and symmetric structures, and arranging block-diagonal structures. One of the most common approaches to solving NMPC is direct methods, which transform NMPC into a Non-Linear Program (NLP) with the complexity of $O(N^{3})$, where $N$ is the problem size [200]. Some direct methods, such as direct multiple shooting and direct collocation, result in sparse NLPs, whose computation complexity can be reduced to $O(N)$ [201]. Another approach to solve NMPC is single-shooting methods, such as DDP [202] and its variant, Iterative Linear Quadratic Regulator (iLQR) [203], which only retains the first-order derivative approximation of dynamics and exhibits a linear increase in computation over the horizon [204]. With proper exploitation of the sparsity structure through the hypergraph approach, [205] shows improvement of the nonlinear solver in computation efficiency. Recently, a numerically robust solver, fatrop [206], solves constrained OCP problems in a direct multiple-shooting fashion efficiently by employing a structure-exploiting linear solver. In NMPC problems, fatrop achieves comparable solve time compared to acados SQP solver while retaining similar numerical robustness to the interior-point-method-based ipopt solver. Furthermore, AdaptiveNLP leverages the previous NLP structure to significantly reduce the overhead and update time for constructing the current NLP [207]. Due to the static nature of most inequality constraints on humanoid robots, such as joint states, actuation, and control barriers function, providing a smaller set of inequality constraints with this memory-aware and adaptive solver has the potential to accelerate humanoid loco-manipulation NMPC. As an increasingly popular approach, structure exploitation has vast potential due to the minimal trade-off in the numerical robustness of the solution.
@@
NMPC Speed-up via Linearization: Another way to tackle computational burdens brought by NMPC is successive linearization, which involves linearization at every timestep around the nominal system state and control input. The linearized dynamics become piece-wise affine, which can be formulated in a large, sparse Quadratic Program (QP) and can be solved online [208, 209]. Aiming for highly efficient optimization-based planning, ReLU-QP [210], a GPU-accelerated QP solver, has improved MPC real-time control frequency in humanoid balancing tasks from original $65$ Hz to more than $1300$ Hz. Successive linearization, however, sacrifices model fidelity and inevitably results in motion errors compared to using accurate nonlinear models. In practice, trading model accuracy for speed is often a preferred strategy because a controller may not be able to track the accurate full-order trajectory with high precision, and therefore it is not practically beneficial to pursue accurate trajectories generated from the full-order, nonlinear model.
@@
NMPC Speed-up via Warm Start: The real-time requirement motivates many researchers to seek a better initialization. One simple yet effective approach is to warm start with the solution from the previous iteration. While promising in reducing computation burden, this approach highly depends on the quality of previous solutions and is sensitive to changes in dynamics or task constraints across time steps, which are common for contact-rich, multi-task loco-manipulation. Another common approach is to offload the computation burden from online to offline, e.g., the gait library [211]. It can be regarded as a specific type of warm-start technique and requires only a cheap online interpolation among the gaits to obtain an approximately optimal full-body trajectory. Similarly, [190] use memory of motions to warm-start an MPC and overcome the sensitivity of initial conditions. A proper initialization from memory usually takes only a few iterations to achieve convergence, which enables online NMPC with full-body dynamics. Combining offline memories with online planning is a promising research direction. However, the key challenge lies in the management of massive trajectories with limited storage. In Sec. VII-E, we discuss a solution from the learning community: learning compact models to distill from large-scale offline trajectories.
@@ V-E1 Interaction with a static environment
Multi-contact MPC for humanoid robots can be solved by optimization-based methods [188, 189]. Among these methods, Centroidal Dynamics (CD) is the most common model due to its accurate representation of contact forces and the system’s centroidal momentum. Despite the model’s accuracy, CD contain a nonlinear term derived from the cross-product between the state (CoM position) and control (contact wrench), posing a challenge to trajectory optimization. Using multi-contact MPC as a motion planning technique also has limited dynamic locomotion capabilities because it treats arms and legs uniformly as general contacts, making it less efficient at handling frequent contact switches compared to pure locomotion models such as the linear inverted pendulum model (LIPM). Although MPC has the ability to plan contact with any surface of the robot, the regularization of the planned contact forces usually requires accurate joint torque sensing or whole-body tactile sensing (Sec. III-C), which still has significant space to explore and presents great potential for rich and safe environment interaction.
@@ V-E2 Modeling interaction with a manipulated object
In the context of loco-manipulation MPC for humanoid robotics, modeling strategies for manipulated objects represent a crucial aspect and an area of ongoing research, alongside contact planning. An object can be a free-floating body (e.g. a box), a fixed-base articulated mechanism (e.g. a door or drawer), or actuated joints (e.g. another robot) [221], as shown in Fig. 8 (c-g). Unlike interactions with a static environment, the contact force exerted by an object depends not only on the robot’s joint torque but also on the object’s mass and inertia. As a result, interacting with objects in loco-manipulation tasks introduces significant complexity. Planning such tasks typically requires accurate knowledge of the object’s state and physical properties, especially when handling heavy, irregular, or dynamically moving objects.
To overcome the challenge of unknown object states and properties, adaptive control schemes and online estimation techniques have been proposed to improve robustness, compensating for dynamic effects and varying external loads. For instance, [222] compensate for the dynamic effects of an object as residual dynamics, thereby avoiding the need for extensive predefined object parameters. [223] estimate the robot’s reflected inertia online to compensate for constantly changing external loads. [224] estimate the object’s mass to select the optimal whole-body manipulation strategy for bulky objects. In addition, to address object state feedback, wide-angle camera dense tracking is proposed to aid in tracking large objects [84]. However, application to MPC-based approaches poses additional challenges in terms of prediction of compensation for the object dynamics in the preview horizons, prediction of object motion evolution, and increased density of the online computation load due to such integrations. [225] makes the initial step toward estimating and simplifying the evolution of centroidal momentum through supervised learning to maintain a convex CD-based MPC formulation in humanoid locomotion.
@@
The first approach models the manipulated object as external wrenches and plans the control input to compensate for them [70, 2, 189]. This approach offers a flexible solution, as it integrates well with the MPC regardless of the linearity of the MPC model, which treats all contacts as external wrenches. However, the contact wrench needs to be predefined, for example, to compensate for the gravity of the object or to exert a user-specified pushing force. Obtaining accurate contact wrenches for dynamic tasks like swinging a baseball bat is already inherently challenging, especially when considering their evolution over the entire prediction horizon in MPC. Static/quasi-static assumptions are usually made to neglect the dynamics of the object, resulting in less dynamic loco-manipulation motions. Another aspect to note, the contact wrenches can be applied at the contact location [82] or at the CoM of the robot [28, 2]. In the former setting, the object affects the contact wrenches for both self-balance and object manipulation. In the latter setting, the object affects only the contact wrenches that are responsible for balance, and contact wrenches for object manipulation require additional regulation. For example, the loco-manipulation MPC approach in [2] adjusts the foot contact wrenches to the weight of an object applied to the robot CoM and additionally regulates the object position via hand contact wrenches with a separate controller. Unlike MCP, such loco-manipulation MPC prioritizes mobility over manipulation and typically employs specialized locomotion models, such as a linear inverted pendulum model (LIPM). These models introduce additional assumptions for bipedal locomotion, such as assigning foot contact for locomotion and hand contact for manipulation, maintaining body height, and conserving angular momentum, making them computationally efficient in an online MPC setting but less general.
@@ V-E3 Interaction with a dynamic environment or deformable objects
Dynamically changing environments, such as those with a moving surface [226, 227] or with physical human interactions [28], introduce additional challenges to loco-manipulation planning and control. Similar to a dynamic object manipulation problem, the interaction model between the robot and the dynamic environment is also time-varying. Although one can infuse the dynamics of the object with the robot model to form unified dynamics, it is impractical to model the dynamics of the environment numerically in most cases. Therefore, in an MPC setting, the planner may require sensor feedback to predict the movement of the environment and replan the loco-manipulation motion adaptively [226, 228]. For example, interacting with an environment that involves humans requires the anticipation of human intentions for collaborative manipulation such as lifting payload [28]; see a more challenging recent achievement in direct human-humanoid physical interaction [229]. For such tasks, the interaction force is an important way of communicating intentions, which can be measured as force feedback signals to trigger robot movements. However, the evolution of such sensed forces can not be well-predicted beyond the current timestep for MPC to leverage, suggesting further static/quasi-static assumptions are required. Otherwise, the robot can only treat the dynamic environment as a disturbance and counteract it through reactive control (e.g., whole-body control). Given the challenge of dealing with the changing environment, loco-manipulation in a dynamic environment is largely unexplored.
@@
In addition to rigid objects with regular geometries, deformable objects are ubiquitous in our real world, such as those in caregiving or housekeeping scenarios. Modeling the dynamics of these objects requires a deep understanding of their physical properties and behaviors, such as flexibility, elasticity, and deformation under force. Consequently, simplifications tailored to specific problems and applications are often necessary [230, 231]. For example, to plan the manipulation of a deformable belt, [232] simplify the motion of the belt by representing only its tail movement in a 2D plane. However, to fully exploit the object’s deforming property for effective loco-manipulation, integrating accurate deformable objects [233] into robot models is essential. Although this area is relatively underexplored for humanoid loco-manipulation, such integration opens up significant opportunities beyond basic pick-and-place operations, enabling robots to tackle more intricate and delicate tasks.
@@
Conclusion: With the advanced capabilities of gradient-based numerical optimization in motion planning, MPC is gaining popularity in humanoid loco-manipulation, showcasing numerous variations in recent years of literature. The essence of this method lies in making reasonable choices regarding the dynamics model, constraint, task definitions, and computation requirements. These choices often require expert design and tuning to trade-offs among task versatility, solution feasibility, and optimality. By identifying the computation intensities and proper dynamics representation of the loco-manipulation tasks, one can offset the computation load by introducing simplified models and relaxed constraints in MPC. In addition, the MPC efficiency can greatly benefit from proper solver choices, an evolving area presenting opportunities for research on both solver-level and problem-formulation-level innovations. For further reading, we recommend the survey on MPC for legged and humanoid robots [13]. Additionally, loco-manipulation tasks present further challenges due to the complexity of dynamic interactions with both the environment and objects, which leaves open questions on how to choose and formulate the interactive dynamics effectively based on the specific task requirements in an MPC setting.

### v2

@@ IV-D Challenges in Multi-contact Planning
Key Takeaways: While significant progress has been made in humanoid multi-contact planning, future work should focus on developing more integrated approaches that combine the strengths of search-based, optimization-based, and learning-based methods. Specifically, addressing the computational complexity of CIP and improving real-time performance will be key. Future directions could explore hybrid approaches that incorporate efficient contact sequence generation/contact dynamics, apply contact-implicit constraints in real-time, and achieve learning-based contact prediction to enhance robustness and adaptability in complex loco-manipulation tasks. The readers are recommended to further read the survey on humanoid multi-contact planning [143].
@@ V-A Simplified Models
TABLE III: Recent MPC Approaches on Humanoid Loco-manipulation
Paper Year Robot Model Interaction Modeling Method Locomotion (L) and Manipulation (M) Planning MPC Frequency Solving Method
[183] 2014 SRBM Optimizing external wrench(es) at contact(s) Unified $20$ Hz QP
[2] 2023 SRBM Predefined external force Unified $300$ Hz QP
[184] 2024 SRBM Negligible object dynamics Separated: L: MPC; M: Keyframe interpolation -- QP
[185] 2019 LIPM Negligible object dynamics Separated: L: MPC; M: Teleoperation and retargeting -- QP
[130] 2016 CD Optimizing external wrench(es) at contact(s) Unified offline SQP
[33] 2019 CD Predefined external wrench Unified -- QP
[186] 2014 CD Optimizing external wrench(es) at contact(s) Unified $5$ Hz/offline SQP
[187] 2020 CD Optimizing external wrench(es) at contact(s) Unified $10$ Hz Interior-point
[181] 2023 CD Estimated as external wrench through sensors Unified $5$ Hz Interior-point
[188] 2021 WBD Optimizing external wrench(es) at contact(s) Unified $100$ Hz DDP
@@
On the other hand, the Linear Inverted Pendulum Model (LIPM) has been a popular choice as a linearized dynamics model for humanoid locomotion [27] and in multi-contact [189]. Extending LIPM to loco-manipulation tasks is achieved through teleoperation [185]. However, such a model inherently lacks the ability to address contact interactions and loco-manipulation dynamics, necessitating a lower-level whole-body control for balancing and manipulation tasks.
@@ V-B Whole-body Models
While simplified dynamics models offer computation efficiency, the simplification assumptions cause a detriment to the model’s accuracy and limit the capability of whole-body motion planning. Conversely, whole-body models are more accurate representations of the robot dynamics and are better suited for planning versatile motions and interactions with objects and the environment. NMPC comes into play when constraints or cost functions become nonlinear, for example, dynamics constraints formed by kinodynamics and Whole-Body Dynamics (WBD).
@@
Leveraging joint-space WBD (eqn. (5)) in MPC has gained popularity, particularly for free-floating articulated robots like humanoids. While WBD allows flexible contact modeling, its high nonlinearity and nonconvexity impose significant computational burdens on WBD-based Nonlinear Programs (NLP) or NMPC, making real-time planning challenging. This issue is especially critical for high-DOF humanoids in loco-manipulation tasks, such as payload transport, which require additional modeling of object dynamics and safety constraints. Thus, this section surveys methods to accelerate NMPC while preserving solution accuracy.
@@ V-C Mixed-fidelity Models
Cascaded-fidelity models (a.k.a. hierarchical dynamics) use different models to govern segments of the horizon [191, 192, 193]. These methods typically employ high-fidelity (e.g., full-order) models in near horizons and low-fidelity (e.g., simple) models for distant horizons, thus maintaining the solution accuracy in near horizons while solving the myopic issue by allowing a longer horizon using simple models. This approach could be suitable in loco-manipulation tasks as it would either simplify interaction dynamics as simple external forces or impose the object dynamics as part of CD in the far horizons to allow improved real-time computation compared to full dynamics models.
Another method is to overlap different dynamics models across their horizons. In such cases, achieving a consensus between these overlapped models is necessary. To solve problems with such mixed-fidelity models, [194] decomposes a single TO that incorporates both dynamics into two subproblems and then alternates between the two to achieve consensus. Similarly, [195] alternates between the CD and WBK subproblems. Overall, model simplification over MPC horizons will remain an effective approach [9]. On the other hand, mixed-fidelity models demonstrate superior capability but require careful consideration of combined models.
TABLE IV: NMPC Speed-up Strategies:
Accuracy, Efficiency, and Limitations
Method Accuracy Efficiency Limitation and Challenge
i. Dynamics Modeling
Simplified model (Sec. V-A) low high Possible infeasible motion trajectory due to simplification
Whole-body model (Sec. V-B) very high very low Requires very accurate dynamics modeling of robot and object
Mixed-fidelity model (Sec. V-C) medium medium Motion trajectory discrepancy due to model mismatch
ii. NLP Structure Exploitation
Backward recursion:
 DDP [196] medium high
 iLQR [197] low very high Poor constraint-handling
Derivative calculation:
 Auto-Diff [198] very high low
 Hypergraph [199] high medium May lead to high memory usage
 Exploring sparsity:
 FATROP [200] very high high Requires specific problem structure
 AdaptiveNLP [201] high high Less effective in small-scale NLPs
iii. Linearization
 SQP [202] high medium Scalability to high-order
Successive Linearization [203] medium high Requires good initial guess
 QP Speed-up:
 Condensed [204] very high medium No state trajectory output
 ReLU-QP [205] high very high Only time-invariant states matrices
iv. Warm Start
 Gait Library [206] high very high Constrained by onboard storage
Memory of Motion [188] high low Requires large dataset coverage
Performance scale (low to high): very $\rightarrow$ $\rightarrow$ $\rightarrow$ high $\rightarrow$ very high.
@@ V-D NMPC Speed-up
In addition to tailoring dynamics modeling strategies for specific loco-manipulation setups, solving NMPC can be accelerated using several key methods. Below, we summarize these methods and provide Table LABEL:tab:MPC_speedup to compare their solution accuracy, real-time efficiency, and application limitations.
@@
NMPC Speed-up via Structure Exploitation: NMPC problems often involve complex dynamics and constraints that can be computationally intensive to solve. Exploiting the structure within these problems can significantly enhance their solvability and efficiency, such as extracting variables that directly interact with each other, identifying repetitive and symmetric structures, and arranging block-diagonal structures. One of the most common approaches to solving NMPC is direct methods, which transform the NMPC into a NLP with the complexity of $O(N^{3})$, where $N$ is the problem size [207]. Some direct methods, such as direct multiple shooting and direct collocation, result in sparse NLPs, whose computation complexity can be reduced to $O(N)$ [208]. Another approach to solve NMPC is single-shooting methods, such as DDP [196] and its variant, Iterative Linear Quadratic Regulator (iLQR) [197], which only retains the first-order derivative approximation of dynamics and exhibits a linear increase in computation over the horizon [209]. With proper exploitation of the sparsity structure through the hypergraph approach, [199] shows improvement of the nonlinear solver in computation efficiency. Furthermore, AdaptiveNLP leverages the previous NLP structure to significantly reduce the overhead and update time for constructing the current NLP [201], which is particularly suitable for humanoid robot NMPCs with static kinematics and actuation constraints.
@@
NMPC Speed-up via Linearization: Another way to tackle the computational burdens of NMPC is through successive linearization, which involves linearization at every timestep around the nominal system state and control input. The linearized dynamics become piece-wise affine, which can be formulated in a large, sparse Quadratic Program (QP) and can be solved online [203, 210]. SQP solves NLP problems by iteratively linearizing them into QPs and computing search directions at each step. While breaking down an NLP into QPs is viable, it requires efficient and scalable QP solvers. Aiming for high efficiency, ReLU-QP [205], a GPU-accelerated QP solver, has improved linear convex MPC real-time control frequency in high dimensional space balancing tasks from the original 206 Hz to up to 2600 Hz. While providing reasonably efficient real-time deployment, linearization techniques highly rely on linearization assumptions, which compromise model fidelity (accuracy), leading to motion errors compared to full nonlinear models. However, trading accuracy for speed is often preferred in humanoid robot control, as controllers may struggle to precisely track full-order trajectories by NMPC, making highly accurate solutions not practically beneficial.
@@
NMPC Speed-up via Warm Start: The real-time requirement motivates many researchers to seek a more effective initialization. One simple yet effective approach is to warm start with the solution from the previous iteration. Another common approach is to offload the computation burden from online to offline, e.g., the gait library [206]. It can be regarded as a specific type of warm-start technique and requires only a cheap online interpolation among the gaits to obtain an approximately optimal full-body trajectory. Similarly, [188] uses memory of motions to warm-start an MPC and overcome the sensitivity of initial conditions. However, the key challenge lies in the management of massive trajectories with limited storage. In Sec. VII-E, we discuss a solution from the learning community: learning compact models to distill from large-scale offline trajectories.
@@ V-E1 Interaction with a static environment
Multi-contact MPC for humanoid robots can be solved by optimization-based methods [186, 187]. Among these methods, CD is the most common model due to its accurate representation of contact forces and the system’s centroidal momentum. Despite the model’s accuracy, CD contains a nonlinear term derived from the cross-product between the state (CoM position) and control (contact wrench), posing a challenge to trajectory optimization. Using multi-contact MPC as a motion planning technique also has limited dynamic locomotion capabilities because it treats arms and legs uniformly as general contacts, rendering it less efficient at handling frequent contact switches compared to pure locomotion models such as the LIPM. Although MPC has the ability to plan contact with any surface of the robot, the regularization of the planned contact forces usually requires accurate joint torque sensing or whole-body tactile sensing (Sec. III-C), which still has significant space to explore and presents great potential for rich and safe environment interaction.
@@ V-E2 Modeling interaction with a manipulated object
In the context of loco-manipulation MPC for humanoid robotics, modeling strategies for manipulated objects represent a crucial aspect and an area of ongoing research, alongside contact planning. An object can be a free-floating body (e.g. a box), a fixed-base articulated mechanism (e.g. a door or drawer), or actuated joints (e.g. another robot) [219], as shown in Fig. 8 (c-g). Unlike interactions with a static environment, the contact force exerted by an object depends not only on the robot’s joint torque but also on the object’s mass and inertia. As a result, interacting with objects in loco-manipulation tasks introduces significant complexity. Planning such tasks typically requires accurate knowledge of the object’s state and physical properties, especially when handling heavy, irregular, or dynamically moving objects. As a result, loco-manipulation tasks become significantly more complex, often requiring precise knowledge of an object’s state and physical properties—especially when dealing with heavy, irregular, or dynamic objects.
To address unknown object states and properties, adaptive control and online estimation techniques have been proposed to enhance robustness against dynamic effects and external loads. For example, [220] compensates for residual dynamics without requiring predefined object parameters, while [221] estimates reflected inertia to handle changing loads. Nozawa et al. [222] estimates object mass to optimize manipulation strategies for bulky items. Chappellet et al. [93] leverages wide-angle camera tracking to aid in large object handling. However, integrating these methods with MPC-based approaches introduces challenges, such as predicting object dynamics over preview horizons and managing increased computational load. Bang et al. [223] takes a step toward simplifying centroidal momentum evolution through supervised learning, preserving a convex CD-based MPC formulation for humanoid locomotion.
@@
The first approach models the manipulated object as external wrenches and plans the control input to compensate for them [79, 2, 187]. This approach offers a flexible solution, as it integrates well with the MPC regardless of the linearity of the model, which treats all contacts as external wrenches. However, the contact wrench needs to be predefined, for example, to compensate for the gravity of the object or to exert a user-specified pushing force. Obtaining accurate contact wrenches for dynamic tasks like swinging a baseball bat is already inherently challenging, especially when considering their evolution over the entire prediction horizon in MPC. Static/quasi-static assumptions are usually made to neglect the dynamics of the object, resulting in less dynamic loco-manipulation motions. Another aspect to note is that contact wrenches can be applied at the contact location [91] or at the robot’s CoM [33, 2]. In the former setting, the object affects the contact wrenches for both self-balance and object manipulation. In the latter setting, the object affects only the contact wrenches that are responsible for balance, and contact wrenches for object manipulation require additional regulation. For example, the loco-manipulation MPC approach in [2] adjusts the foot contact wrenches to the weight of an object applied to the robot CoM and additionally regulates the object position via hand contact wrenches with a separate controller. Unlike MCP, such loco-manipulation MPC prioritizes mobility over manipulation and typically employs specialized locomotion models, such as a linear inverted pendulum model (LIPM). These models introduce additional assumptions for bipedal locomotion, such as assigning foot contact for locomotion and hand contact for manipulation, maintaining body height, and conserving angular momentum, making them computationally efficient in an online MPC setting albeit being less general.
@@ V-E3 Interaction with a dynamic environment or deformable objects
Dynamically changing environments, such as those with a moving surface [224] or with physical human interactions [33], introduce additional challenges to loco-manipulation planning and control. Similar to a dynamic object manipulation problem, the interaction model between the robot and the dynamic environment is also time-varying. Although one can infuse the dynamics of the object with the robot model to form unified dynamics, it is impractical to model the dynamics of the environment numerically in most cases. Therefore, in an MPC setting, the planner may require sensor feedback to predict the movement of the environment and replan the loco-manipulation motion adaptively [224]. For example, interacting with an environment that involves humans requires the anticipation of human intentions for collaborative manipulation such as lifting payload [33]; see a more challenging recent achievement in direct human-humanoid physical interaction [225]. For such tasks, the interaction force is an important way of communicating intentions, which can be measured as force feedback signals to trigger robot movements. However, the evolution of such sensed forces can not be well-predicted beyond the current timestep for MPC to leverage, suggesting further static/quasi-static assumptions are required. Otherwise, the robot can only treat the dynamic environment as a disturbance and counteract it through reactive control (e.g., whole-body control). Given the challenge of dealing with the changing environment, loco-manipulation in a dynamic environment is largely unexplored.
@@
In addition to rigid objects with regular geometries, deformable objects are ubiquitous in our real world, such as those in caregiving or housekeeping scenarios. Modeling the dynamics of these objects requires a deep understanding of their physical properties and behaviors, such as flexibility, elasticity, and deformation under force. Consequently, simplifications tailored to specific problems and applications are often necessary [226, 227]. For example, to plan the manipulation of a deformable belt, [228] simplifies the motion of the belt by representing only its tail movement in a 2D plane. However, to fully exploit the object’s deforming property for effective loco-manipulation, integrating accurate deformable objects [229] into robot models is essential. Although this area is relatively underexplored for humanoid loco-manipulation, such integration opens up significant opportunities beyond basic pick-and-place operations, enabling robots to tackle more intricate and delicate tasks.
@@
Key Takeaways for Model Predictive Control: With the advanced capabilities of gradient-based numerical optimization in motion planning, MPC is gaining popularity in humanoid loco-manipulation, showcasing numerous variations in recent years of literature. The essence of this method lies in making reasonable choices regarding the dynamics model, constraint, task definitions, and computation requirements. These choices often require expert design and tuning to trade-offs among task versatility, solution feasibility, and optimality. By identifying the computation intensities and proper dynamics representation of the loco-manipulation tasks, one can offset the computation load by introducing simplified models and relaxed constraints in MPC. In addition, the MPC efficiency can greatly benefit from proper solver choices, an evolving area presenting opportunities for research on both solver-level and problem-formulation-level innovations. For further reading, we recommend the survey on MPC for legged and humanoid robots [14]. Additionally, loco-manipulation tasks present further challenges due to the complexity of dynamic interactions with both the environment and objects, which leaves open questions on how to choose and formulate the interactive dynamics effectively based on the specific task requirements in an MPC setting.